cart at a constant velocity for a distance of 20.0 m. She pushes in a direction 24.0o below the h A 32.0-N frictional force opposes the motion of the cart. (a) What is the magnitude of the force that the shopper exerts? (b) Determine the work done by the pushing force (c) Determine the work done by the frictional force (d) Determine the work done by the gravitational force 6 P008
Answer #1

P = Pushing Force

t = angle below horizontal

F = friction force

Resolving Horizontally :-

P * cos(t) = F

a) P = F / cos(t)

= 32.0 / cos(24)

= 32.0 / 0.9135

= 35.03 N

b) Work done by P

= P * Cos(t) * 20

= F * 20

= 32 * 20

= 640 J

c) Work done by F

= - F * 20

= -32 * 20

= -640 J

d) Work done by gravitational force is "0"

since there is no displacement in the vertical direction.